1) If a 20.0 L tank filled with propane has a temperature of 35.0 degrees Celsius and a pressure of 1100.0kPa., the number of moles of propane in the cylinder is ______mol. *


2) The volume occupied by one mole of hydrogen iodide gas is at a pressure of 855 kPa and a temperature of 28.0 degrees Celsius. How many grams of hydrogen iodide does this vessel contain? *


1
Expert's answer
2020-06-29T07:52:18-0400

1)  PV = nRT

Here,   P = 1100.0 kPa,  T = 308.15 K,  V = 20.0 L,  R = 8.314 kPa L K-1 mol-1

 n = PV/RT = (1100.0×20.0) / (8.314×308.15) mol

                 = 8.587 mol  (Answer)

There is 8.578 moles propane in the cylinder.

2)    PV = nRT  

Here, n= 1, P = 855 kPa,  T = 301.15 K,  R = 8.314 kPa L K-1 mol-1

             V = (1×8.314×301.15) / 855 L

                = 2.928 L

   (P1V1)/ T1 = (P2V2) / T2

According to question here, P1= 855 kPa,  T1 = 301.15 K,   V1 = 2.928 L

                                             P2= 101.325 kPa,  T2 = 293.15 K  [ use NTP values ]

                     V2 = (855×2.928×293.15) / (301.15×101.325) L

                            = 24.05 L

At NTP,   22.4 L hydrogen iodide = 1 mole hydrogen iodide

               24.05 L hydrogen iodide = 24.05/22.4 mole  

                                                           =   1.07  mole

Molecular weight of hydrogen iodide is 127.911 g/mol

1 mole hydrogen iodide = 127.911 g

1.07 mole hydrogen iodide = 136.86 g (Answer)

136.86 grams of hydrogen iodide contain this vessel.